(1) C&A's potato chip filling process has a lower specification limit of 9.5 oz and an upper specification limit of 10.5 oz. The standard deviation is 0.3 oz. and the mean is 10 oz. What is the process capability index (Cp) for the chip filling process?

(a) 0.56
(b) 1.11
(c) 0.33
(d) 3.33


(2) Which of the following represents customer requirement from a process?

(a) Sample size in control charts
(b) USL and LSL
(c) Process standard deviation
(d) UCL and LCL


(3) The mean of a process is 50, and the standard deviation is 2. The company uses statistical control process (SPC) to monitor the process. It uses sample size on 4, and 3 sigma (z=3) control limits. What are the Upper Control Limit (UCL) and Lower Control Limit (LCL)?

(a) LCL=48 and UCL=52
(b) LCL=44 and UCL=56
(c) LCL=47 and UCL=53
(d) LCL=49 and UCL=51


(4) If a company narrows the control limits for a process without making any other changes, which of the following will happen?

(a) The process will generate more products not acceptable by customers
(b) The process will not be stopped even when assignable cause is present
(c) The capability index for the process will reduce to a lower value
(d) The process will be stopped more often suspecting assignable cause
Answer #1

1. process capabilityindex = (USL - LSL)/6Sigma = 1/6(0.3)= 0.56 answer (a) 0.56

2. ans. (d) UCL and LCL

3. UCL = mean +3sigma = 50 +3(2) = 56. LCL = mean - 3sigma = 50 - 3(2) = 44 ans. (b) LCL=44 and UCL=56

4.ans. (d) The process will be stopped more often suspecting assignable cause